### v1

# Origin of Bright Quantum Emissions with High Debye-Waller factor in Silicon Nitride

Abstract
Silicon nitride has emerged as a promising photonic platform for integrated single-photon sources, yet the microscopic origin of the recently observed bright quantum emissions remains unclear. Using hybrid density functional theory, we show that the negatively charged NV center (NV-) in the C1h configuration exhibits a linearly polarized zero-phonon line (ZPL) at 2.46 eV, with a radiative lifetime of 9.01 ns and a high Debye-Waller (DW) factor of 33%. We further find that the C1h configuration is prone to a pseudo-Jahn-Teller distortion, yielding two symmetrically equivalent defect structures that emit bright, linearly polarized ZPL at 1.80 eV with a lifetime of 10.17 ns and an increased DW factor of 41%. These nitrogen-vacancy-related defects explain the origins of visible quantum emissions, paving the way for deterministic and monolithically integrated silicon-nitride quantum photonics.
Bright single-photon emitters operating near the visible spectrum have attracted significant interest, exemplified by the 2 eV emitter in hBN [1] and the NV- center in diamond [2]. Hybrid integration approaches, which combine different materials for the emission source and photonic elements, increase fabrication complexity and introduce coupling losses at interfaces [1, 3]. A unified photonic platform with built-in single-photon emitters is therefore crucial for achieving direct, scalable, and low-loss integration of quantum light sources. Silicon nitride has emerged as a state-of-the-art quantum photonic platform, offering a wide bandgap suitable for hosting quantum emitters. Recently, room-temperature single-photon emissions near 2 eV have been observed in silicon nitride [4, 3]. These emitters are bright, stable, linearly polarized, and exhibit high purity. Most importantly, monolithic integration of these emitters within the silicon nitride photonic platform has been demonstrated [5]. However, the microscopic origin of these emitters remains unidentified.
Senichev et al. [4] reported intrinsic single-photon emitters in silicon nitride, emitting in the range of 568-722 nm (2.18-1.72 eV) at room temperature. Using a similar synthesis procedure, Martin et al. [3] observed several emitters with ZPL emission around 548 nm (2.26 eV) with a 7 nm deviation at 4.2 K. In both studies, silicon nitride films were grown under nitrogen-rich conditions on SiO2/Si substrates. Analysis of 130 emitters revealed an average second-order autocorrelation value of 0.2 at zero delay, confirming high single-photon purity. Fluorescence lifetime measurements under pulsed excitation for 44 emitters yielded lifetimes ranging from 1.4 to 3.9 ns, with an average of 2.4 ns. The emitters exhibited bright and stable emission without noticeable blinking or bleaching for up to 100 s at saturation power. More recently, Chen et al. [6] identified single-photon emitters with emission between 570 and 582 nm (2.18-2.13 eV), showing large Debye-Waller factors (50-74%) and lifetimes of 3-8 ns. These emitters were attributed to defects in the silica substrate, despite being synthesized using the same method as in Refs. [4, 3]. Given these consistent optical signatures yet differing interpretations, identifying the microscopic nature of these quantum emitters is essential for enabling targeted quantum-optics measurements and deterministic defect creation in integrated Si3N4 photonic platforms.
Here, we identify the microscopic origin of the 2 eV single-photon emitters observed in silicon nitride. Using hybrid density-functional theory, we investigate the thermodynamic stability, electronic structure, and optical properties of the nitrogen-vacancy-antisite (NV, or NV center) defect in -Si3N4. We show that this defect exhibits a bright, linearly polarized optical emission whose ZPL, radiative lifetime, and DW factor are in close agreement with experimentally reported emitter characteristics. We further reveal that a pseudo-Jahn-Teller distortion creates additional symmetry-broken configurations that remain optically active and fall within the same spectral window.
-Si3N4 crystallizes in the centrosymmetric P63/m space group [7], with two inequivalent nitrogen sites (C3h and ) and a single silicon site of symmetry. In this work, we consider the nitrogen vacancy at the C3h site (V) and the associated nitrogen-antisite-vacancy complex (NV). Defects are modeled in a 280-atom hexagonal supercell with -point sampling of the brillouin zone. Calculations are performed using the hybrid Heyd-Scuseria-Ernzerhof (HSE) functional [8] and projector augmented-wave potentials [9], as implemented in VASP [10]. A Hartree-Fock exchange fraction of yields a band gap of 6.1 eV, in good agreement with experiment [1] [see Sec. S1 of the Supplemental Material (SM) [12]; see also Refs. [1, 2, 3, 4, 5, 17, 7, 8, 9, 10] therein]. Unit-cell optimizations used a 449 Monkhorst-Pack k-mesh [22] and a 400 eV plane-wave cutoff, giving relaxed lattice constants Å and Å, consistent with Ref. [23]. Supercell geometries were relaxed using PBE functional [24] until forces were below 0.001 eV/Å, followed by electronic structure calculation using hybrid-functional. Excited-state energies and geometries were obtained using the SCF method [25] and spin polarization is explicitly considered in all the calculations. Electron-phonon coupling was analyzed using PHONOPY [26] and DEFECTPL [27, 7]. Finite-size corrections for charged defects were applied using the eFNV scheme [3] via DOPED [4]. Further computational details are provided in Sec. S2 of SM [12].
From defect formation energy calculations, we find that the nitrogen vacancy (V) is the most stable intrinsic defect in -Si3N4, consistent with previous computational and experimental studies [28, 29]. As shown in Fig. 1(a) and Fig. S1 in the SM [12], V exhibits the lowest formation energy under both nitrogen-rich and silicon-rich conditions. The nitrogen-rich growth environments employed in recent experiments [4, 3, 6] also favor the formation of stable N defects. Furthermore, thermal annealing used to activate the emitters [4] can promote migration of N and V, enabling NV center formation throughout the Si3N4 matrix. Unlike the NV center in diamond, which constitutes an extrinsic substitutional nitrogen adjacent to a carbon vacancy, the NV defect in -Si3N4 originates from an intrinsic nitrogen antisite (N) adjacent to V [Fig. 1(b)]. As shown in Fig. 1(a), the NV center is stable in charge states +1, 0, and 1 within the bandgap as the Fermi level varies between the valence band maximum (VBM) and conduction band minimum (CBM). The neutral NV center (NV0) has a singlet ground state, whereas the charged NV+ and NV- defects exhibit doublet ground states, each hosting a single unpaired electron. In contrast, the V is paramagnetic with one unpaired electron, while its positive and negative charge states exhibit singlet ground states.
The relaxed geometry of V exhibits C3h point-group symmetry, in contrast to the C3v symmetry reported by Nanataki et al. [30]. In the neutral charge state, the defect hosts a single unpaired electron occupying a pair of nearly degenerate majority-spin orbitals, which drives a strong Jahn-Teller (JT) distortion of the local environment. In the high-symmetry configuration, the three neighboring Si atoms form an almost perfect equilateral triangle with C3h symmetry. Upon JT distortion, however, this triangle becomes scalene, and the symmetry is lowered to C1h. The resulting symmetry breaking lifts the degeneracy of the defect orbitals and produces a large splitting of approximately 2.9 eV (see Fig. S2 in the SM [12]). In contrast, the positively and negatively charged V states do not exhibit a JT instability and therefore preserve C3h symmetry. When a nitrogen atom substitutes a neighboring Si site, forming the NV complex, structural relaxation yields a C1h configuration. This structure contains a mirror plane passing through the substituted N atom and the two adjacent Si atoms.
Nitrogen vacancy introduces three planar dangling bonds derived from the 3s and 3p orbitals of the neighboring Si atoms. Their interaction produces localized defect states within the bandgap (see Fig. S2 in the SM [12]). When a nitrogen atom substitutes a nearby Si site to form the NV center, hybridization between the substituted N and surrounding N atoms alters the vacancy levels. Additionally, the defect can capture an electron, leading to a localized state in the minority-spin channel just above the VBM [see C1h configuration in Fig. 2(a)]. The highest occupied and lowest unoccupied minority-spin defect states are separated by 3.94 eV. In the NV- charge state, these two localized states transform as irreducible representations, denoted as (occupied) and (unoccupied). Optical transitions between them are linearly polarized in the -plane, producing a ZPL at 2.46 eV with a transition dipole moment (see Eq. S16 in SM [12]) of 4.7 Debye. This is in close agreement with the experimentally reported ZPL near 2.26 eV [4, 3]. The corresponding radiative rate and lifetime (see Eq. S15 in SM [12]) are 111 MHz and 9.01 ns, respectively. Our calculated values also agree well with ZPL (2.15 eV) and lifetime (38 ns) reported by Chen et al. [6]. In the majority-spin channel [Fig. 2 (a)], the occupied and unoccupied states give rise to a second dipole-allowed transition polarized along , as only transforms as . This transition yields a ZPL at 0.85 eV, within the telecommunication spectral range, with a transition dipole of 7.2 Debye and a radiative rate and lifetime of 10.5 MHz and 95.24 ns, respectively.
The NV- center in the C1h configuration exhibits an imaginary phonon mode. This mode corresponds to a localized out-of-plane oscillation of the substituted nitrogen (NSi) between neighboring nitrogen atoms [Fig. 2(b)], indicating that the C1h structure is metastable. It originates from the shorter N-N bond compared to the Si-N bond. A small displacement along this mode drives the system toward a lower-symmetry configuration, in which the substituted nitrogen moves closer to one neighboring nitrogen, breaking the mirror symmetry. Since the C1h configuration is nondegenerate yet still distorts, this behavior is classified as a pseudo-Jahn-Teller (PJT) effect [31]. We denote the resulting configuration as C1h-PJT [Fig. 2(b)]. The two possible distortions, toward either neighboring nitrogen, are energetically degenerate, with approximately 1 eV lower energy than the high-symmetry C1h structure. In realistic environments, particularly near interfaces or under local strain, the NV- center may adopt any configuration near the symmetric C1h and distorted C1h-PJT geometries. The distortion of the C1h configuration of the NV- defect modifies the energy and symmetry of the localized states [see Fig. 2(a)], altering its optical properties. In the minority-spin channel, C1h-PJT configuration exhibits a ZPL at 1.80 eV and a radiative lifetime of 10.17 ns, in close agreement with experimental reports [4]. Given the variability in the local environment of NV- centers in real samples, their optical properties are expected to fall close to that of the C1h and C1h-PJT configurations.
Next, we analyzed the electron-phonon coupling of this transition using the method of Alkauskas et al. [17]. The photoluminescence spectra show three phonon sidebands, also reflected in the calculated electron-phonon spectral function [Fig. 3]. The 87 meV phonon mode with the largest partial Huang-Rhys (HR) factor is quasilocalized with a localization ratio of 14.4, while the spectral peaks arise from multiple phonon modes with different degrees of localization [see Fig. 3 and Fig. S3 in SM [12]]. The first peak near 55 meV agrees with the first sideband in the range 38.6-57.4 meV observed in experiment [6], with additional peaks at 83 and 118 meV. The total HR factor is 0.9, corresponding to a DW factor of 41, which is close to the experimentally measured range of 50-70 reported by Chen et al. [6]. These DW values are significantly higher than the DW factor of 3 for the well-known NV- center in diamond, even though the ZPL energy and excited-state lifetime are comparable, at approximately eV and 12 ns [32, 33]. The moderate electron-phonon interaction (HR factor = 0.9) ensures that a large fraction of the optical emission remains in the zero-phonon line [see Fig. 3(b)]. This behavior arises because the structural relaxation between the ground and excited states is relatively small, with a change in configuration coordinate [Q, see Eq. S11 in SM [12]] of only 0.36 amu1/2Å. For the C1h configuration, the HR and DW factors are 1.11 and 33, respectively. The strength of the electron-phonon coupling, expressed by the HR factor, is reflected in the calculated Q for both the C1h (0.44 amu1/2Å) and C1h-PJT (0.36 amu1/2Å) geometries. Earlier studies [4, 3] proposed that the observed photoluminescence features might originate from a single family of defects with a complex structure. Our results support this interpretation and indicate that the experimentally observed emission is consistent with nitrogen-vacancy-related centers in silicon nitride, rather than defects in a silica host [6]. Further controlled synthesis and targeted experiments will be essential to verify the microscopic origin of these single-photon emitters. Our findings will guide future efforts to isolate and engineer bright quantum emitters in silicon nitride.
In summary, we identify nitrogen-vacancy-related defects, particularly NV-, as the likely origin of the recently observed 2 eV single-photon emission in silicon nitride. The NV- exhibits a linearly polarized ZPL emission at 2.46 eV in the C1h configuration, with a lifetime of 9.01 ns and a high DW factor of 33%. We further show that pseudo-Jahn-Teller distortions generate symmetrically equivalent configurations (C1h-PJT), which modify the electronic structure and emission characteristics with a ZPL, lifetime, and DW factor of 1.80 eV, 10.17 ns, and 41%, respectively. The moderate electron-phonon coupling during electronic transition gives higher quantum efficiency. Our study provides a microscopic understanding of bright quantum emitters and suggests that controlled creation of nitrogen-vacancy-related defects could enable deterministic integration of single-photon sources in silicon-nitride photonic circuits. This work lays the foundation for the design and implementation of scalable quantum photonic devices based on silicon nitride.
The authors thank the Materials Research Centre (MRC), the Supercomputer Education and Research Centre (SERC), and the Solid State and Structural Chemistry Unit (SSCU) at the Indian Institute of Science, Bangalore, for access to computing resources. This work was supported by the DST-Nanomission program of the Department of Science and Technology, Government of India (Grant No. DST/NM/TUE/QM-1/2019). S.M. acknowledges funding from the PMRF fellowship (ID No. 0201908). The authors also acknowledge support from the Institute of Eminence (IoE) scheme of the Ministry of Human Resource Development, Government of India.
Supplemental Materials: Origin of Bright Quantum Emissions with High Debye-Waller factor in Silicon Nitride
I Section S1: Bulk Properties
The bandgaps of -Si3N4 obtained from various theoretical approaches and experimental measurements are summarized in Table 1. Our calculated bandgap using hybrid functional parameters AEXX = 0.3 and HFSCREEN = 0.2 closely matches both the computed value and the experimentally reported bandgap of 6.5 eV for crystalline -Si3N4, as measured by linear fitting of EELS-STEM data by Asano et al. [1]. However, discrepancies persist among experimentally reported bandgap values in the literature.
| Method | PBE | DDH | HSE06 | HSEa | G0W0b | Exp.c |
| Eg (eV) | 4.4 | 5.8 | 5.8 | 6.1 | 6.2 | 6.5 |
aWith AEXX=0.3 and HFSCREEN=0.2
bGW oneshot calculation
cExperimental bandgap from linear fit in EELS-STEM [1]
II Section S2: Theoretical Methodology
II.0.1 Defect Formation Energy
The defect formation energy of a defect in charge state , as a function of the Fermi level , is calculated using the expression [2, 3]:
Here, and are the total energies of the defect-containing and pristine bulk supercells, respectively. The term accounts for the exchange of atoms with reservoirs, where is the number of atoms of species added () or removed (), and is the corresponding chemical potential.
The chemical potentials are defined as follows:
Under N-rich conditions:
Under Si-rich conditions:
The remaining chemical potentials are derived from the constraint:
where is the total energy per formula unit of crystalline Si3N4.
The term represents the energy required to exchange electrons with the reservoir defined by the Fermi level . includes finite-size corrections for potential alignment and charged defect interactions. We employ the extended Freysoldt–Neugebauer–Van de Walle (eFNV) correction as implemented in the DOPED package [3, 4].
II.0.2 Photoluminescence and Electron-Phonon Coupling
To study the electron-phonon coupling and photoluminescence (PL) spectrum for defect-to-defect transitions, we use the theory of Huang and Rhys [5], as demonstrated by Alkauskas et al. [6] and implemented in custom built DEFECTPL package [7]. The normalized PL lineshape is given by:
where the optical spectral function is defined as:
Here, and are vibrational wavefunctions of the ground and excited states, respectively, and is the energy of the ground-state vibrational level.
Assuming similar phonon modes in both states and using the generating function approach [8, 9], the spectral function becomes:
where is the generating function and determines the ZPL width. The function is the Fourier transform of the electron-phonon coupling spectral function :
The total Huang-Rhys factor is:
Each partial HR factor is given by:
with:
Here, is the mass of atom , and are atomic coordinates in the ground and excited states, and is the component of the normalized displacement vector of the phonon mode.
The spectral function is then:
The total HR factor quantifies the electron-phonon coupling and is reflected in the configuration coordinate change:
where is the difference in atomic coordinates of atom between ground and excited states.
II.0.3 Debye-Waller Factor and Localization Ratio
The Debye-Waller factor represents the fraction of emission in the ZPL, is:
Phonon mode localization is estimated using the localization ratio:
where is the number of atoms in the supercell and is the inverse participation ratio [6]:
II.0.4 Radiative Transition Rate and Lifetime
The radiative transition rate and lifetime are calculated as:
Here, is the refractive index of -Si3N4, and is the magnitude of the transition dipole moment in eÅ. The constants , , and are used in units of , Å/s, and eV·s, respectively.
The transition dipole moment between ground and excited single-particle states is computed using the VaspBandUnfolding package [10]:
III Section S3: Defect Thermodynamics in Si-rich Conditions
Under Si-rich conditions, the nitrogen vacancy is found to be the most stable defect with a maximum formation energy reaching up to 4 eV for the neutral charge state. However, nitrogen substitution at silicon sites becomes energetically unfavorable with higher formation energies, resulting in less stable antisite and antisite-vacancy defects involving nitrogen.
IV Section S4: Electronic Structure of Neutral Nitrogen Vacancy
V Section S5: Phonon Mode Localization

### v2

@@
 Given these consistent optical signatures yet differing interpretations, identifying the microscopic nature of these quantum emitters is essential for enabling targeted quantum-optics measurements and deterministic defect creation in integrated Si3N4 photonic platforms.
-Here, we identify the microscopic origin of the 2 eV single-photon emitters observed in silicon nitride.
+In this letter, we identify the microscopic origin of the 2 eV single-photon emitters observed in silicon nitride.
 Using hybrid density-functional theory, we investigate the thermodynamic stability, electronic structure, and optical properties of the nitrogen-vacancy-antisite (NV, or NV center) defect in -Si3N4.
@@
 A Hartree-Fock exchange fraction of yields a band gap of 6.1 eV, in good agreement with experiment [1] [see Sec.
-S1 of the Supplemental Material (SM) [12]; see also Refs.
+S1 of the Supplementary Information (SI); see also Refs.
 [1, 2, 3, 4, 5, 16, 7, 8, 9, 10] therein].
@@
 [22].
+Furthermore, the calculated formation enthalpy of -8.96 eV per formula unit is in good agreement with that of the experimentally reported value of -8.58 eV [23].
 Supercell geometries were relaxed using PBE functional [24] until forces were below 0.001 eV/Å, followed by electronic structure calculation using hybrid-functional.
 Excited-state energies and geometries were obtained using the SCF method [25] and spin polarization is explicitly considered in all the calculations.
-Electron-phonon coupling was analyzed using PHONOPY [26] and DEFECTPL [27, 7].
+Electron-phonon coupling was analyzed using PHONOPY [26] and DEFECTPL [27, 7, 28].
 Finite-size corrections for charged defects were applied using the eFNV scheme [3] via DOPED [4].
 Further computational details are provided in Sec.
-S2 of SM [12].
+S2 of SI.
 From defect formation energy calculations, we find that the nitrogen vacancy (V) is the most stable intrinsic defect in -Si3N4, consistent with previous computational and experimental studies [29, 30].
@@
 1(a) and Fig.
-S1 in the SM [12], V exhibits the lowest formation energy under both nitrogen-rich and silicon-rich conditions.
+S1 in the SI, V exhibits the lowest formation energy under both nitrogen-rich and silicon-rich conditions.
 The nitrogen-rich growth environments employed in recent experiments [4, 3, 6] also favor the formation of stable N defects.
 Furthermore, thermal annealing used to activate the emitters [4] can promote migration of N and V, enabling NV center formation throughout the Si3N4 matrix.
+Our nudged elastic band calculations suggest that the formation of the NV defect from V involves a high migration barrier of 2.66 eV, as shown in Fig.
+S4 of the SI.
 Unlike the NV center in diamond, which constitutes an extrinsic substitutional nitrogen adjacent to a carbon vacancy, the NV defect in -Si3N4 originates from an intrinsic nitrogen antisite (N) adjacent to V [Fig.
@@
 As shown in Fig.
-1(a), the NV center is stable in charge states +1, 0, and 1 within the bandgap as the Fermi level varies between the valence band maximum (VBM) and conduction band minimum (CBM).
+1(a), the NV center is stable in charge states 2 to 2 within the bandgap as the Fermi level varies between the valence band maximum (VBM) and conduction band minimum (CBM).
 The neutral NV center (NV0) has a singlet ground state, whereas the charged NV+ and NV- defects exhibit doublet ground states, each hosting a single unpaired electron.
@@
 The resulting symmetry breaking lifts the degeneracy of the defect orbitals and produces a large splitting of approximately 2.9 eV (see Fig.
-S2 in the SM [12]).
+S2 in the SI).
 In contrast, the positively and negatively charged V states do not exhibit a JT instability and therefore preserve C3h symmetry.
@@
 Nitrogen vacancy introduces three planar dangling bonds derived from the 3s and 3p orbitals of the neighboring Si atoms.
-Their interaction produces localized defect states within the bandgap (see Fig.
-S2 in the SM [12]).
-When a nitrogen atom substitutes a nearby Si site to form the NV center, hybridization between the substituted N and surrounding N atoms alters the vacancy levels.
-Additionally, the defect can capture an electron, leading to a localized state in the minority-spin channel just above the VBM [see C1h configuration in Fig.
+Their interaction produces localized Kohn-Sham (KS) levels within the bandgap (see Fig.
+S2 in the SI).
+When a nitrogen atom substitutes a nearby Si site to form the NV center, hybridization between the substituted N and surrounding N atoms alters the KS levels.
+The negatively charged defect features an occupied KS level just above the VBM and an unoccupied level located 3.94 eV higher in energy [see C1h configuration in Fig.
 2(a)].
-The highest occupied and lowest unoccupied minority-spin defect states are separated by 3.94 eV.
 In the NV- charge state, these two localized states transform as irreducible representations, denoted as (occupied) and (unoccupied).
+Both the ground and excited states are spin doublets and represented as .
 Optical transitions between them are linearly polarized in the -plane, producing a ZPL at 2.46 eV with a transition dipole moment (see Eq.
-S16 in SM [12]) of 4.7 Debye.
+S16 in SI) of 4.7 Debye.
 This is in close agreement with the experimentally reported ZPL near 2.26 eV [4, 3].
 The corresponding radiative rate and lifetime (see Eq.
-S15 in SM [12]) are 111 MHz and 9.01 ns, respectively.
+S15 in SI) are 111 MHz and 9.01 ns, respectively.
 Our calculated values also agree well with ZPL (2.15 eV) and lifetime (38 ns) reported by Chen et al.
 [6].
+Similar nitrogen antisite-vacancy complex defect has also been observed as source of 2 eV SPE in GaN [32].
 In the majority-spin channel [Fig.
@@
 This mode corresponds to a localized out-of-plane oscillation of the substituted nitrogen (NSi) between neighboring nitrogen atoms [Fig.
-2(b)], indicating that the C1h structure is metastable.
+2(b)], indicating that the C1h configuration is an unstable saddle point in the potential energy surface.
 It originates from the shorter N-N bond compared to the Si-N bond.
-A small displacement along this mode drives the system toward a lower-symmetry configuration, in which the substituted nitrogen moves closer to one neighboring nitrogen, breaking the mirror symmetry.
+A small displacement along this mode drives the system toward a lower-symmetry configuration, in which the substituted nitrogen moves closer to one neighboring nitrogen, breaking the mirror symmetry [see Fig.
+2(b) and Fig.
+S7 in SI].
 Since the C1h configuration is nondegenerate yet still distorts, this behavior is classified as a pseudo-Jahn-Teller (PJT) effect [33].
-We denote the resulting configuration as C1h-PJT [Fig.
+This distortion results in C1 point group symmetry and is denoted as C1h-PJT from now onwards [Fig.
 2(b)].
 The two possible distortions, toward either neighboring nitrogen, are energetically degenerate, with approximately 1 eV lower energy than the high-symmetry C1h structure.
-In realistic environments, particularly near interfaces or under local strain, the NV- center may adopt any configuration near the symmetric C1h and distorted C1h-PJT geometries.
 The distortion of the C1h configuration of the NV- defect modifies the energy and symmetry of the localized states [see Fig.
@@
 In the minority-spin channel, C1h-PJT configuration exhibits a ZPL at 1.80 eV and a radiative lifetime of 10.17 ns, in close agreement with experimental reports [4].
-Given the variability in the local environment of NV- centers in real samples, their optical properties are expected to fall close to that of the C1h and C1h-PJT configurations.
+In realistic samples, the NV- center may experience diverse local environments, such as those near interfaces or regions under strain, that can alter its electronic and optical characteristics.
+Given that the optical response is sensitive to subtle changes in the atomic configuration, as illustrated by the differences in properties of the C1h and C1h-PJT structures, such local variations likely contribute to the spread and variability observed in experimental measurements.
 Next, we analyzed the electron-phonon coupling of this transition using the method of Alkauskas et al.
 [16].
-The photoluminescence spectra show three phonon sidebands, also reflected in the calculated electron-phonon spectral function [Fig.
+The photoluminescence spectra show a phonon sideband with three peaks, also reflected in the calculated electron-phonon spectral function [Fig.
 3].
@@
 3 and Fig.
-S3 in SM [12]].
-The first peak near 55 meV agrees with the first sideband in the range 38.6-57.4 meV observed in experiment [6], with additional peaks at 83 and 118 meV.
+S3 in SI].
+The first peak near 55 meV agrees with the first peak of phonon sideband in the range 38.6-57.4 meV observed in experiment [6], with additional peaks at 83 and 118 meV.
 The total HR factor is 0.9, corresponding to a DW factor of 41, which is close to the experimentally measured range of 50-70 reported by Chen et al.
@@
 This behavior arises because the structural relaxation between the ground and excited states is relatively small, with a change in configuration coordinate [Q, see Eq.
-S11 in SM [12]] of only 0.36 amu1/2Å.
+S11 in SI] of only 0.36 amu1/2Å.
 For the C1h configuration, the HR and DW factors are 1.11 and 33, respectively.
 The strength of the electron-phonon coupling, expressed by the HR factor, is reflected in the calculated Q for both the C1h (0.44 amu1/2Å) and C1h-PJT (0.36 amu1/2Å) geometries.
+We note that PBE functional underestimates the ground-to-excited state configurational displacement Q relative to that calculated from HSE functional [36], leading to reduced HR factor and overestimated Debye-Waller values.
+While SCAN provides an improvement over PBE, it still underestimates experimental HR values for the NV- center in diamond [37].
+Thus, our reported HR values represent a reliable lower bound.
 Earlier studies [4, 3] proposed that the observed photoluminescence features might originate from a single family of defects with a complex structure.
@@
 We further show that pseudo-Jahn-Teller distortions generate symmetrically equivalent configurations (C1h-PJT), which modify the electronic structure and emission characteristics with a ZPL, lifetime, and DW factor of 1.80 eV, 10.17 ns, and 41%, respectively.
-The moderate electron-phonon coupling during electronic transition gives higher quantum efficiency.
+The moderate electron-phonon coupling during the electronic transition gives a higher emission in the ZPL compared to the phonon-side band.
 Our study provides a microscopic understanding of bright quantum emitters and suggests that controlled creation of nitrogen-vacancy-related defects could enable deterministic integration of single-photon sources in silicon-nitride photonic circuits.
@@
 II Section S2: Theoretical Methodology
-II.0.1 Defect Formation Energy
+II.0.1 S2.1: Defect Formation Energy
 The defect formation energy of a defect in charge state , as a function of the Fermi level , is calculated using the expression [2, 3]:
@@
 includes finite-size corrections for potential alignment and charged defect interactions.
-We employ the extended Freysoldt–Neugebauer–Van de Walle (eFNV) correction as implemented in the DOPED package [3, 4].
-II.0.2 Photoluminescence and Electron-Phonon Coupling
+We employ the extended Freysoldt-Neugebauer-Van de Walle (eFNV) correction as implemented in the DOPED package [3, 4].
+II.0.2 S2.2: SCF Calculation
+For the C configuration, the ground and excited states are both (doublets), while for the C-PJT distorted structure (C1 symmetry), those are both .
+The ground and excited states of the C configuration can be represented by single Slater determinants as
+respectively, where the overbar denotes electron occupation in the minority-spin channel.
+Similarly, in the C-PJT configuration, the ground and excited states can be represented by single Slater determinants as
+respectively. Therefore, the SCF approach is applicable here because these correspond to doublet-to-doublet transitions dominated by single Slater determinants.
+II.0.3 S2.3: Photoluminescence and Electron-Phonon Coupling
 To study the electron-phonon coupling and photoluminescence (PL) spectrum for defect-to-defect transitions, we use the theory of Huang and Rhys [5], as demonstrated by Alkauskas et al.
@@
 where is the difference in atomic coordinates of atom between ground and excited states.
-II.0.3 Debye-Waller Factor and Localization Ratio
+II.0.4 S2.4: Debye-Waller Factor and Localization Ratio
 The Debye-Waller factor represents the fraction of emission in the ZPL, is:
@@
 where is the number of atoms in the supercell and is the inverse participation ratio [6]:
-II.0.4 Radiative Transition Rate and Lifetime
+II.0.5 S2.5: Radiative Transition Rate and Lifetime
 The radiative transition rate and lifetime are calculated as:
@@
 IV Section S4: Electronic Structure of Neutral Nitrogen Vacancy
-V Section S5: Phonon Mode Localization
+V Section S5: Localization Ratio of Phonon Modes
+VI Section S6: Migration Barrier
+We performed nudged elastic band (NEB) calculation to evaluate the feasibility of a neighbouring nitrogen atom migrating from the nitrogen site adjacent to a silicon vacancy to the silicon vacancy site.
+As shown in Figure S4, our results indicate that this process involves a substantial energy barrier of 2.66 eV, as shown below.
+Furthermore, once the nitrogen atom is at the silicon site, the reverse transition back to the nitrogen site is also highly unlikely, with a calculated barrier of 2.04 eV.
+VII Section S7: Local Geometry
+The bond lengths between atoms around NSiV defect in the C1h configuration and the C1h-PJT configuration are shown in Figure S5.